OpenStudy (anonymous):

What is the middle term of the product of (x - 4)(x - 3)? -x x -7x

OpenStudy (anonymous):

-7x

mathslover (mathslover):

Note the formula : (x-a)(x-b) = x^2 - (a+b)x + ab Middle term is : -(a+b)x = -(4+3)x = -7x